how many solutions do this linear system have? 2x-6y=-14,-2x+3y=7

Answer by checkley71(8403)

2x-6y=-14
-6y=2x-14
y=2x/-6-14/-6
y=-x/3+7/3 (red line)
-2x+3y=7
3y=2x+7
y=2x/3+7/3 (green line)
(graph 300x200 pixels, x from -6 to 5, y from -10 to 10, of TWO functions y = -x/3 +7/3 and y = 2x/3 +7/3).
ONLY 1 SOLUTION (0,7/3).